OpenStudy (anonymous):

there are 370 people attending a concert. each adult paid $5 admission and each child paid $2. if $1220 was received for admission,, how many adults and how many children attended the concert?

Directrix (directrix):

Let A = the number of adults attending the concert Let C = the number of children attending the concert So, 1) A + C = 370 The amount of money made from adults is 5*A The amount of money made from children is 2*C So, 2) 5A + 2C = 1220

Directrix (directrix):

Your task is to solve the two simultaneous equations for the variables A and C: 1) A + C = 370 2) 5A + 2C = 1220 @harry.g1234